# Generalization and Knowledge Transfer in Abstract Visual Reasoning Models

### v1

@@ Abstract
We study generalization and knowledge reuse capabilities of deep neural networks in the domain of abstract visual reasoning (AVR), employing Raven’s Progressive Matrices (RPMs), a recognized benchmark task for assessing AVR abilities. Two knowledge transfer scenarios referring to the I-RAVEN dataset are investigated. Firstly, inspired by generalization assessment capabilities of the PGM dataset and popularity of I-RAVEN, we introduce Attributeless-I-RAVEN, a benchmark with four generalization regimes that allow to test generalization of abstract rules applied to held-out attributes. Secondly, we construct I-RAVEN-Mesh, a dataset that enriches RPMs with a novel component structure comprising line-based patterns, facilitating assessment of progressive knowledge acquisition in transfer learning setting. The developed benchmarks reveal shortcomings of the contemporary deep learning models, which we partly address with Pathways of Normalized Group Convolution (PoNG) model, a novel neural architecture for solving AVR tasks. PoNG excels in both presented challenges, as well as the standard I-RAVEN and PGM setups.
@@ 1 Introduction
Generalization, the ability of a model to perform well on unseen data, remains a fundamental challenge in deep learning (DL). While DL methods have demonstrated remarkable achievements in various domains, their generalization capabilities are often questioned, particularly in tasks that demand abstract problem-solving and reasoning skills [6]. One such domain is abstract visual reasoning (AVR) [31, 47, 44, 26] that encompasses tasks requiring (human) fluid intelligence – an aspect of human cognition believed to be crucial for reasoning in never-encountered settings [42, 5]. The most popular AVR tasks are Raven’s Progressive Matrices (RPMs) [38, 39], which constitute a common problem found in human IQ tests. Typical RPMs comprise two components – the context panels arranged in a grid with the bottom-right panel missing and up to 8 answer panels, out of which only one correctly completes the matrix. Solving an RPM instance requires identification of underlying abstract rules applied to certain attributes of the objects composing the instance (see Fig. 1 for an illustrative example).
@@
Design of computational methods capable of tackling RPMs has for decades been an active area of research [8, 12, 10, 24, 22, 45]. Consequently, a number of works considered automatic creation of RPM datasets [30, 48, 29] and a wide suite of predictive models [14, 13] were proposed, with DL methods showing the most promising performance [53, 25]. While this rapid progress led to exceeding the human level in particular problem setups [50, 33], a fundamental challenge of generalization to novel problem settings remains largely unattained.
@@
Initial works designed several RPM datasets [30, 48, 16], however, measuring generalization was not their focus. While some works explored knowledge transfer between related tasks [29, 46], the complexity of the datasets was limited and consequently they didn’t pose a challenge for contemporary DL methods. To measure generalization in modern DL models, the PGM dataset was introduced [2]. PGM defines eight generalization regimes, each specifying the distribution of objects, rules and attributes in train and test splits. For instance, in the Held-out Triples split, a given rule–object–attribute triplet (e.g. Progression on Object’s Size) was assigned only to one of the two splits. In effect, the models were tested on triplet combinations different from training ones, allowing to assess their generalization capabilities. A subsequent work proposed RAVEN [54], another RPM dataset with enriched perceptual complexity of matrices instantiated in seven visual configurations (Center, 2x2Grid, 3x3Grid, Left-Right, Up-Down, Out-InCenter, Out-InGrid). Moreover, the benchmark is characterized by a moderate sample size, i.e. 70K instances, compared to 1.42M RPMs per each of the eight regimes in PGM. Due to this size disparity, subsequent research gravitated towards RAVEN and its revised variants (I-RAVEN [17] and RAVEN-Fair [3]), which didn’t require substantial computational resources to train DL models.
@@
Drawing inspiration from the generalization assessment capabilities of PGM and the broad adoption of RAVEN, this paper proposes a novel suite of generalization challenges stemming from I-RAVEN [17] (a revised variant of RAVEN that removes a bias in RAVEN’s answer panels). The proposed suite of benchmarks allows assessing generalization of AVR models without requiring substantial computational resources.
@@
Firstly, we introduce Attributeless-I-RAVEN, comprising four generalization regimes corresponding to specific held-out attributes ({Position, Type, Size, Color}), resp. Each regime includes four rules ({Constant, Progression, Arithmetic, Distribute Three}) and the training matrices adhere to the Constant rule for the respective attribute, whereas test matrices employ a rule different from Constant for this attribute. In effect, each regime comprises different distributions of training and test data. Note that unlike PGM, in which regimes are oriented over sets of attributes or their range of values, each Attributeless-I-RAVEN regime targets generalization to novel rules concerning a specific attribute, enabling concept-based assessment of generalization capabilities [34, 36].
@@
Secondly, we propose I-RAVEN-Mesh, a variant of I-RAVEN with a new grid-like structure overlaid on the matrices. The dataset enables assessing generalization to incrementally added structures and progressive knowledge acquisition in a transfer learning (TL) setting.
In investigations involving nine contemporary AVR DL models, we observed that the introduced benchmarks present a substantial challenge for the tested methods. This prompted the development of Pathways of Normalized Group Convolution (PoNG), a novel AVR model that excels in both problem setups: generalization to held-out attributes and incremental knowledge acquisition.
Our main contributions can be summarized as follows:
@@
We introduce the Attributeless-I-RAVEN (A-I-RAVEN) dataset that enables measuring generalization across four regimes.
@@
We propose a new neural architecture for solving AVR tasks termed PoNG, which excels in addressing both introduced challenges.
@@ 2 Related Work
In recent years, a variety of AVR problems and corresponding datasets have emerged [4, 35, 9, 37, 41, 19, 15, 56] and several attempts have been made to measure generalization in contemporary AVR models based on the introduced benchmarks. In particular, distinct visual configurations were employed in RAVEN to assess how a model trained on one configuration performs on the remaining ones [54, 43, 57]. Although in such a setting the visual aspects of train/test matrices come from different distributions, the underlying rules and attributes remain the same. In contrast, A-I-RAVEN enables studying the generalization of rules applied to held-out attributes, shifting the focus from perception towards reasoning. Besides RPMs, the limits of generalization have been explored in other AVR tasks as well. Visual Analogy Extrapolation Challenge evaluates model’s capacity for extrapolation [49]. However, such specialized datasets might favor models that explicitly embed the notion of extrapolation in their design and aim for being invariant only to specific attributes such as object size or location. Differently, our benchmarks allow verifying the model’s capacity to learn a given concept from the data and generalize it to novel settings. This perspective links our work to the recent literature on concept learning [36, 34]. However, the concept-oriented benchmarks that originate from ARC [6] remain largely unsolved by DL models and pose a significant challenge even for leading multi-modal large language models [32]. In contrast, both benchmarks proposed in this work are attainable by DL models, though further advances in generalization abilities of the models are necessary to consider them solved.
@@
Preliminary attempts to solve RPMs with DL models involve WReN [2] that reasons over object relations using Relation Network [40], or SRAN [17] that relies on a hierarchical architecture with panel encoders devoted to particular image groups. A common theme enabling generalization in DL models is to explicitly identify RPM objects. To this end, RelBase [43] employs Attend-Infer-Repeat [7], an unsupervised scene decomposition method, STSN [33] utilizes Slot attention [23] to decompose matrix to slots containing particular objects and Temporal Context Normalization (TCN) [49] to normalize latent matrix panel representations in a task-specific context, and MRNet [3] presents a multi-scale architecture. SCL [50] proposes the scattering transformation, CoPINet [55] and CPCNet [52] rely on contrastive architectures, PredRNet [51] learns to minimize the prediction error, and SCAR [28] adapts its computation to the structure of the considered matrix. Despite the high variety of AVR models, experiments on the introduced benchmarks reveal their shortcomings in terms of generalization and knowledge transfer. In this context, we propose PoNG, a new AVR model that excels in the presented tasks by combining parallel architecture, weight sharing, and tactical normalization.
3 Methods
@@
The set of attributes in I-RAVEN is Position, Number, Type, Size, Color and the set of rules is Constant, Progression, Arithmetic, Distribute Three. For an attribute and a dataset split , where train, validation, test, we define the set of rules applicable to in split by . In I-RAVEN all rule–attribute pairs are valid in all splits:
3.1 Attributeless-I-RAVEN
@@
To probe generalization in DL models, we present A-I-RAVEN, a benchmark composed of 4 generalization regimes. Example matrices are illustrated in Fig. 2, with additional samples provided in Appendix A. Each regime, denoted as A/<Attribute> (e.g., A/Type), corresponds to a single held-out attribute. In train and validation splits, all matrices have the Constant rule applied to this attribute. In the test split, the held-out attribute is governed by a rule sampled from . In effect, during training, the model doesn’t see rule–attribute combinations required to solve test matrices. There are no rule-related constraints on the remaining attributes. Since Position and Number attributes are tightly coupled (e.g., it’s impossible to increase object numerosity while keeping their position constant), we allocate a single generalization regime, A/Position, to cover both attributes. In summary, for a held-out attribute , we have:
@@
DT | Three line counts are sampled and spread among images in a given row. | |
@@
DT | Three line arrangements are sampled and spread among images in a given row. |
@@ 3.2 I-RAVEN-Mesh
The other of the proposed benchmarks is designed to probe progressive knowledge acquisition in a TL setting. I-RAVEN-Mesh extends I-RAVEN by introducing a novel visual component overlaid on top of the existing I-RAVEN components (see Fig. 3). Though the dataset can serve as a learning challenge on its own, the main motivation behind its introduction is to employ models pre-trained on I-RAVEN and fine-tune them on I-RAVEN-Mesh with a configurable train sample size, facilitating analysis of their TL performance. The mesh grid comprises from 1 to 12 lines placed in predefined locations. The set of available lines covers the inner and outer edges of a grid (12 lines in total). The mesh component has two attributes: , which govern the count and location of lines, respectively. To each attribute a rule can be applied. Table 1 describes the effect of applying a given rule–attribute pair to the mesh component.
To generate the mesh component of an I-RAVEN-Mesh matrix, we sample one of the two attributes and a corresponding rule that governs its values. As the attributes often depend on each other (e.g., it’s impossible to increase the number of lines while keeping their position constant), we don’t constrain the value of the other attribute. The rule–attribute pairs for the base I-RAVEN components are generated in the same way as in the original dataset. To generate answers to the matrix, we follow the impartial algorithm proposed in I-RAVEN [17]. In addition, each matrix contains at least one incorrect answer that differs from the correct one only in the mesh component, ensuring that the solver has to identify the correct rule governing the mesh component in order to solve the matrix. To facilitate training with an auxiliary loss, in which the model additionally predicts the representation of rules governing the matrix [2], we extend the base set of rule annotations with ones concerning the Mesh component.
3.3 Pathways of Normalized Group Convolution (PoNG)
In initial experiments, we’ve found out that SOTA AVR models struggle in the proposed generalization challenges. Consequently, we introduce PoNG, a novel model that outcompetes baselines across a number of problem settings. The model follows a typical two-stage design. Firstly, it generates an embedding of each image panel. Then, it aggregates representations of matrix panels to predict the index of the correct answer. PoNG is illustrated in Fig. 4 and its details are described in Appendix E.
Let denote an RPM, where is the set of image panels comprising context panels and answer panels , is a greyscale image of height and width , is the one-hot encoded index of the correct answer, is the multi-hot encoded representation of matrix rules of dimensionality using sparse encoding [27]. In each experiment , while is determined by the number of matrix components in the corresponding dataset ( for I-RAVEN-Mesh, otherwise; see Appendix D for details).
Panel encoder.
The first component of the model has the form , where is the input panel embedding of dimensionality . Following RelBase [43], the module comprises blocks of the same architecture. Each block includes parallel pathways that build high-level and low-level features, resp. The first one contains convolutional blocks, each with 2D convolution, ReLU, and Batch Normalization (BN) [18]. The second one contains 2D max pooling followed by 2D convolution. The sum of both pathway results forms the block output. Differently from RelBase, we flatten the height and width dimensions of the resultant embedding, pass it through a linear layer with ReLU, flatten the channel and spatial dimensions, and pass the tensor through a feed-forward residual block with Layer Normalization (LN) [1]. Finally, we concatenate the tensor with a position embedding (a learned -dimensional vector for each cell in the context grid), leading to .
Reasoner.
The second component of the model has the form , where is the panel embedding of ’th answer. For each answer panel, the reasoner produces embedding that describes how well the considered answer fits into the matrix context. Panel embeddings are stacked and processed by a sequence of reasoning blocks interleaved with bottleneck layers for dimensionality reduction. Each reasoning block comprises BN and parallel pathways, outputs of which are added together to form the output of the block. Next, the latent representation is passed through adaptive average pooling, flattened, processed with a linear layer with ReLU, passed through BN and projected with a linear layer to .
Pathways.
The key aspect of the reasoner module are its pathways. Each takes an input tensor of shape , where is the batch size, is the number of channels, and is the feature dimension. In the first reasoning block and corresponds to the number of panel embeddings in the considered group. Pathways are described as follows: P1 – a pointwise 1D convolution layer that mixes panel features at each spatial location; P2 – a sequence of 2 blocks, each comprising 1D convolution, ReLU, and BN, that builds higher level features spanning neighbouring spatial locations; P3 – analogous to P2, but 1D convolution is replaced with a group 1D convolution that splits the tensor into several groups along the channel dimension, applies a 1D convolution with shared weights to each group, and adds together the representations of each group; P4 – analogous to P3, but groups are arranged into pairs concatenated along the channel dimension and processed with a 1D convolution with shared weights. In contrast to [21], the proposed group convolution layers in both P3 and P4 apply TCN [49] to the outputs in each group. In the first layer P3 and P4 split the input tensor into 3 groups, which allows for producing embeddings of each matrix row and each pair of rows, resp. Though we apply the pathways block in the RPM context, we envisage it as a generic module, also applicable to other settings involving a set of vector representations of shape .
Answer prediction.
Eight representations of the context matrix filled-in with the respective answer, , are processed with three prediction heads. The target head employs two linear layers interleaved with ReLU to produce score describing how well the answer aligns with the matrix context. The aggregate rule head computes the sum of inputs and processes it with two linear layers interleaved with ReLU, producing a latent prediction of matrix rules . The target-conditioned rule head processes its input through a linear layer and computes a weighted sum of the resultant embeddings with weights given by the predicted probability distribution over the set of possible answers , where denotes softmax. The model is trained with a joint loss function , where denotes sigmoid, CE cross-entropy, BCE binary cross-entropy, and are balancing coefficients.
4 Experiments
We assess generalization of state-of-the-art models for solving RPMs on A-I-RAVEN, evaluate progressive knowledge acquisition on I-RAVEN-Mesh, and conduct an ablation study to showcase the contribution of the respective modules that constitute PoNG.
Experimental setup.
In all experiments we use the Adam optimizer [20] with , , and a batch size set to . Learning rate is initialized to and reduced -fold (at most 3 times) if no progress is seen in the validation loss in 5 subsequent epochs, and training stops early in the case of 10 epochs without progress. Unless stated otherwise, each model configuration was trained 3 times with a different seed, and we report mean and standard deviation for these runs. In each experiment, we utilize training, validation, and test matrices, following the standard data split protocol taken in prior works [54, 17]. Experiments were run on a worker with a single NVIDIA DGX A100 GPU.
Baselines.
In addition to the simple CNN-LSTM baseline [2], we assess generalization of SOTA AVR models including WReN [2], CoPINet [55], RelBase [43], SCL [50], SRAN [17], CPCNet [52], PredRNet [51], and STSN [33]. For direct comparison, we evaluate all models on I-RAVEN following the above-described experimental setup.
Reproducibility.
To guarantee reproducibility of experiments, we use a fixed set of random seeds and turn off hardware and framework features concerning indeterministic computation wherever possible. Together with the code, we provide the full training script that can be used to run all training jobs. The training job is packaged as a Docker image with fixed dependencies to isolate the configuration of the training environment. The released code allows for generation of all datasets from scratch, eliminating the dependency on file-hosting services required to distribute the data. The code for reproducing all experiments is publicly accessible at: https://github.com/mikomel/raven.
I-RAVEN | Mesh | A/Color | A/Pos. | A/Size | A/Type | |
@@ 4 Experiments
WReN | ||||||
PoNG (ours) | 95.9 | 89.3 | 80.3 | 79.3 | 73.5 | 59.4 |
Generalization on Attributeless-I-RAVEN.
In the first set of experiments we evaluate all considered models on four generalization regimes of A-I-RAVEN. The results are presented in Table 2, along with the reference results on I-RAVEN and I-RAVEN-Mesh. PoNG outperforms all selected baselines across all settings. Among baseline models, the best results on A/Color are achieved by RelBase, followed by SCL. In the remaining attributeless regimes, SCL outperforms other baselines with RelBase taking the second place. Interestingly, the top 3 models rely on rather shallow architectures, yet outcompete other methods that rely on a deeper layout, such as SRAN or STSN. This suggests that parameter-efficient AVR models not only excel in solving RPMs but also generalize better.
Generalization regimes of A-I-RAVEN pose a bigger challenge than the base dataset. While PoNG, the best performing model, achieved test accuracy on I-RAVEN, on A-I-RAVEN regimes it scored from (on A/Type) to (on A/Color). Fig. 5 displays the difference in PoNG’s performance on test and validation splits. On I-RAVEN and I-RAVEN-Mesh the difference is negligible, as in these datasets both splits follow the same distribution. However, the difference in attributeless regimes is significant, which indicates the need for further research on generalization. We conclude that A-I-RAVEN provides a suite of challenging regimes of variable complexity, in which even the best performing models are far from solving all test matrices.

### v2

# A-I-RAVEN and I-RAVEN-Mesh: Two New Benchmarks for Abstract Visual Reasoning

@@ Abstract
We study generalization and knowledge reuse capabilities of deep neural networks in the domain of abstract visual reasoning (AVR), employing Raven’s Progressive Matrices (RPMs), a recognized benchmark task for assessing AVR abilities. Two knowledge transfer scenarios referring to the I-RAVEN dataset are investigated. Firstly, inspired by generalization assessment capabilities of the PGM dataset and popularity of I-RAVEN, we introduce Attributeless-I-RAVEN (A-I-RAVEN), a benchmark with generalization regimes that allow to systematically test generalization of abstract rules applied to held-out attributes at various levels of complexity (primary and extended regimes). In contrast to PGM, A-I-RAVEN features compositionality, a variety of figure configurations, and does not require substantial computational resources. Secondly, we construct I-RAVEN-Mesh, a dataset that enriches RPMs with a novel component structure comprising line-based patterns, facilitating assessment of progressive knowledge acquisition in transfer learning setting. We evaluate 13 strong models from the AVR literature on the introduced datasets, revealing their specific shortcomings in generalization and knowledge transfer.
@@ 1 Introduction
Generalization, the ability of a model to perform well on unseen data, remains a fundamental challenge in deep learning (DL). While DL methods have demonstrated remarkable achievements in various domains, their generalization capabilities are often questioned, particularly in tasks that demand abstract problem-solving and reasoning skills Chollet (2019). One such domain is abstract visual reasoning (AVR) Mitchell (2021); Stabinger et al. (2021); van der Maas et al. (2021); Małkiński and Mańdziuk (2023) that encompasses tasks requiring (human) fluid intelligence – an aspect of human cognition believed to be crucial for reasoning in never-encountered settings Carpenter et al. (1990). The most popular AVR tasks are Raven’s Progressive Matrices (RPMs) Raven (1936); Raven and Court (1998), which constitute a common problem found in human IQ tests. Typical RPMs comprise two components – the context panels arranged in a grid with the bottom-right panel missing and up to 8 answer panels, out of which only one correctly completes the matrix. Solving an RPM instance requires identification of underlying abstract rules applied to certain attributes of the objects composing the instance (see Fig. 1 for an illustrative example).
@@
Design of computational methods capable of tackling RPMs has for decades been an active area of research Evans (1964); Foundalis (2006); Lovett et al. (2007); Kunda et al. (2010). Consequently, a number of works considered automatic creation of RPM datasets Matzen et al. (2010); Wang and Su (2015); Mańdziuk and Żychowski (2019) and a wide suite of predictive models Hernández-Orallo et al. (2016); Hernández-Orallo (2017) were proposed, with DL methods showing the most promising performance Yang et al. (2022); Małkiński and Mańdziuk (2025). While this rapid progress led to exceeding the human level in particular problem setups Wu et al. (2020); Mondal et al. (2023), a fundamental challenge of generalization to novel problem settings remains largely unattained.
@@
Initial works designed several RPM datasets Matzen et al. (2010); Wang and Su (2015); Hoshen and Werman (2017), however, measuring generalization was not their focus. While some works explored knowledge transfer between related tasks Mańdziuk and Żychowski (2019); Tomaszewska et al. (2022), the complexity of the datasets was limited and consequently they didn’t pose a challenge for contemporary DL methods. To measure generalization in modern DL models, the PGM dataset was introduced Barrett et al. (2018). PGM defines eight generalization regimes, each specifying the distribution of objects, rules and attributes in train and test splits. For instance, in the Held-out Triples split, a given rule–object–attribute triplet (e.g. Progression on Object’s Size) was assigned only to one of the two splits. In effect, the models were tested on triplet combinations different from training ones, allowing to assess their generalization capabilities. A subsequent work proposed RAVEN Zhang et al. (2019a), another RPM dataset with enriched perceptual complexity of matrices instantiated in seven visual configurations (Center, 2x2Grid, 3x3Grid, Left-Right, Up-Down, Out-InCenter, Out-InGrid). Moreover, the benchmark is characterized by a moderate sample size, i.e. instances, compared to RPMs per each of the eight regimes in PGM. Due to this size disparity, subsequent research gravitated towards RAVEN and its revised variants (I-RAVEN Hu et al. (2021) and RAVEN-Fair Benny et al. (2021)), which didn’t require substantial computational resources to train DL models.
Contribution.
@@
Drawing inspiration from the broad adoption of RAVEN and the generalization assessment capabilities of PGM, this paper proposes a novel suite of generalization challenges stemming from I-RAVEN Hu et al. (2021) (a revised variant of RAVEN that removes a bias in RAVEN’s answer panels). However, unlike I-RAVEN, the proposed suite of benchmarks allows for a direct assessment of the generalization and knowledge transfer of AVR models. Compared to PGM, our datasets feature compositionality and variety of figure configurations, and their processing doesn’t require substantial computational resources. Furthermore, they include structural annotations, which are utilized, for example, in recent neuro-symbolic approaches Zhang et al. (2021, 2022).
@@
First, we introduce Attributeless-I-RAVEN (A-I-RAVEN), comprising 10 generalization regimes. The primary regimes correspond to specific held-out attributes ({Position, Type, Size, Color}), resp. The training matrices in these regimes adhere to the Constant rule for the respective attribute, whereas test matrices employ a rule different from Constant for this attribute (i.e., Progression, Arithmetic, or Distribute Three). Moreover, we propose extended regimes: of them feature a held-out attribute pair, while another replace the Constant rule in the training set with each remaining rule. In effect, each regime comprises different distributions of training and test data.
@@
Next, we propose I-RAVEN-Mesh, a variant of I-RAVEN with a new grid-like structure overlaid on the matrices. The dataset enables assessing generalization to incrementally added structures and progressive knowledge acquisition in a transfer learning (TL) setting.
Investigations involving contemporary AVR DL models reveal that the introduced benchmarks present a substantial challenge for the tested methods, raising the need for further advancements in this area.
The key contributions of the paper are summarized below.
@@
We introduce the A-I-RAVEN dataset that enables measuring generalization across regimes.
@@ 2 Related Work
In recent years, a variety of AVR problems and corresponding datasets have emerged Nie et al. (2020); Fleuret et al. (2011); Qi et al. (2021); Shanahan et al. (2020); Hill et al. (2019); Zhang et al. (2020) and several attempts have been made to measure generalization in contemporary AVR models based on the introduced benchmarks. In particular, distinct visual configurations were employed in RAVEN to assess how a model trained on one configuration performs on the remaining ones Zhang et al. (2019a); Spratley et al. (2020); Zhuo and Kankanhalli (2021). Although in such a setting the visual aspects of train/test matrices come from different distributions, the underlying rules and attributes remain the same. In contrast, A-I-RAVEN enables studying the generalization of rules applied to held-out attributes, shifting the focus from perception towards reasoning. Besides RPMs, the limits of generalization have been explored in other AVR tasks as well. Visual Analogy Extrapolation Challenge evaluates model’s capacity for extrapolation Webb et al. (2020). However, such specialized datasets might favor models that explicitly embed the notion of extrapolation in their design and aim for being invariant only to specific attributes such as object size or location. Differently, our benchmarks allow verifying the model’s capacity to learn a given concept from the data and generalize it to novel settings. This perspective links our work to the recent literature on concept learning Moskvichev et al. (2023). However, the concept-oriented benchmarks that originate from ARC Chollet (2019) remain largely unsolved by DL models and pose a significant challenge even for leading multi-modal large language models Mitchell et al. (2023). In contrast, both benchmarks proposed in this work are attainable by DL models, though further advances in generalization abilities of the models are necessary to consider them solved.
@@
Preliminary attempts to solve RPMs with DL models involve WReN Barrett et al. (2018) that reasons over object relations using Relation Network Santoro et al. (2017), or SRAN Hu et al. (2021) that relies on a hierarchical architecture with panel encoders devoted to particular image groups. A common theme enabling generalization in DL models is to explicitly identify RPM objects. To this end, RelBase Spratley et al. (2020) employs Attend-Infer-Repeat, an unsupervised scene decomposition method, STSN Mondal et al. (2023) utilizes Slot attention Locatello et al. (2020) to decompose matrix to slots containing particular objects and Temporal Context Normalization (TCN) Webb et al. (2020) to normalize latent matrix panel representations in a task-specific context, DRNet Zhao et al. (2024) relies on a dual-stream design, and MRNet Benny et al. (2021) presents a multi-scale architecture. SCL Wu et al. (2020) proposes the scattering transformation, CoPINet Zhang et al. (2019b) and CPCNet Yang et al. (2023b) rely on contrastive architectures, PredRNet Yang et al. (2023a) learns to minimize the prediction error, ALANS Zhang et al. (2021) and PrAE Zhang et al. (2022) employ neuro-symbolic architectures, and SCAR Małkiński and Mańdziuk (2024b) adapts its computation to the structure of the considered matrix. Despite the high variety of AVR models, experiments on the introduced benchmarks reveal their shortcomings in terms of generalization and knowledge transfer.
3 Proposed datasets
@@
The set of attributes in I-RAVEN is Position, Number, Type, Size, Color and the set of rules is Constant, Progression, Arithmetic, Distribute Three. For attribute and a dataset split , where train, val., test, we define the set of rules applicable to in split by . In I-RAVEN all rule–attribute pairs are valid in all splits:
3.1 Attributeless-I-RAVEN (A-I-RAVEN)
@@
To probe generalization in DL models, we present A-I-RAVEN, a benchmark composed of generalization regimes. Example matrices are illustrated in Fig. 2, with additional samples provided in Appendix A. Each regime defines a set of held-out attributes , each with a corresponding rule . In train and validation splits, held-out attribute is governed by . In the test split, is governed by a different rule sampled from . In effect, during training, the model doesn’t see rule–attribute combinations required to solve test matrices. There are no rule-related constraints on the remaining attributes. In summary, we have:
We define primary regimes with that correspond to individual held-out attributes (), denoted as A/<Attribute> (e.g., A/Type). Since Position and Number attributes are tightly coupled (e.g., it’s impossible to increase cardinality of objects while keeping their position constant), we allocate a single generalization regime, A/Position, to cover both attributes. In addition, we define extended regimes as supplementary generalization challenges. In the first group a pair of attributes is held-out in the training set, i.e. . Specifically, we introduce new regimes: A/ColorSize, A/ColorType, and A/SizeType, based on the respective attribute pairs. In the second group, Constant rule in is replaced with each of the remaining rules, leading to A/Color-Progression, A/Color-Arithmetic, and A/Color-DistributeThree regimes. While this modification could be applied to all the described regimes, we focus on the Color attribute due to its broad range of possible values.
@@
| Distribute Three | Three line counts are sampled and spread among images in a given row. | |
@@
| Distribute Three | Three line arrangements are sampled and spread among images in a given row. |
@@ 3.2 I-RAVEN-Mesh
The other of the proposed benchmarks is designed to probe progressive knowledge acquisition in a TL setting. I-RAVEN-Mesh extends I-RAVEN by introducing a novel visual component overlaid on top of the existing I-RAVEN components (see Fig. 3). Though the dataset can serve as a learning challenge on its own, the main motivation behind its introduction is to employ models pre-trained on I-RAVEN and fine-tune them on I-RAVEN-Mesh with a configurable train sample size, facilitating analysis of their TL performance. The mesh grid comprises from to lines placed in predefined locations. The set of available lines covers the inner and outer edges of a grid ( lines in total). The mesh component has two attributes: , which govern the count and location of lines, respectively. To each attribute a rule can be applied. Table 1 describes the effect of applying a given rule–attribute pair to the mesh component. To generate the mesh component of an I-RAVEN-Mesh matrix, we sample one of the two attributes and a corresponding rule that governs its values. As the attributes often depend on each other (e.g., it’s impossible to increase the number of lines while keeping their position constant), we don’t constrain the value of the other attribute. The rule–attribute pairs for the base I-RAVEN components are generated in the same way as in the original dataset. To generate answers to the matrix, we follow the impartial algorithm proposed in I-RAVEN Hu et al. (2021). In addition, each matrix contains at least one incorrect answer that differs from the correct one only in the mesh component, ensuring that the solver has to identify the correct rule governing the mesh component in order to solve the matrix. To facilitate training with an auxiliary loss, in which the model additionally predicts the representation of rules governing the matrix Barrett et al. (2018), we extend the base set of rule annotations with ones concerning the Mesh component.
| I-RAVEN† | I-RAVEN (ours) | Mesh | A/Color | A/Position | A/Size | A/Type | |
| ALANS | |||||||
| CPCNet | 98.5 | ||||||
| CNN-LSTM | |||||||
| CoPINet | |||||||
| DRNet | 90.9 | 70.0 | 77.5 | ||||
| MRNet | |||||||
| PrAE | |||||||
| PredRNet | |||||||
| RelBase | 84.9 | ||||||
| SCL | 65.6 | 49.5 | |||||
| SRAN | |||||||
| STSN | |||||||
| WReN |
| A/ColorSize | A/ColorType | A/SizeType | A/Color-P | A/Color-A | A/Color-D3 | |
| ALANS | ||||||
@@
| WReN |
4 Experiments